OpenStudy (anonymous):

0 + 5 = 5 A. commutative property of multiplication B. identity property of addition C. commutative property of addition D. identity property of multiplication

OpenStudy (muzzack):

Obviously we eliminate A and D because it is adding not multiplying

OpenStudy (anonymous):

B?

OpenStudy (muzzack):

Yes, Correct do you know why

OpenStudy (anonymous):

not really...

OpenStudy (muzzack):

Identity property of addition states that the sum of zero and any number or variable is the number or variable itself.

OpenStudy (muzzack):

just keep that in mind

OpenStudy (anonymous):

wait what?

OpenStudy (muzzack):

\(\color{darkgreen}{\huge\``Welcome~to~Open~Study"!!}\)

OpenStudy (muzzack):

Identity property of addition states that the sum of zero and any number or variable is the number or variable itself. For example, 4 + 0 = 4, - 11 + 0 = - 11, y + 0 = y are few examples illustrating the identity property of addition.

OpenStudy (anonymous):

okay thank you

OpenStudy (muzzack):

\(\color{darkgreen}{\huge\ Your~Welcome~!!~}\)